The radius of a quarter circle is 1 millimeter. What is the quarter circle's perimeter?


Use 3.14 for ​𝜋.


1
Expert's answer
2022-06-12T15:31:55-0400

Solution:

"r=1mm;"

Find: "p" = perimeter of quarter circle.

"p=2r+l;" here "l" = arc length.

"l=\frac{2\pi r}{4}=\frac{2\times3.14\times1mm}{4}=1.57mm;"

"p=2\times1mm+1.57mm=3.57mm;"

Answer:

"p=3.57mm."